OpenStudy (anonymous):

.7312 for a present value of $1 means that a certain rate of interest for a certain period of time it will be equal to what

OpenStudy (jamesj):

It means PV of $1 at some particular time in the future = $0.7312 PV of $1 at time t at a discount rate of r is given by \[ PV($1) = \frac{1}{(1+r)^t} \]

OpenStudy (anonymous):

ok but whats the answer

OpenStudy (jamesj):

I thought I answered your question. Ask it again in the best English you have, or even better, copy it from the book.

OpenStudy (anonymous):

i need to know if it will equal $1, over 1$, less than $1, or never equal

OpenStudy (jamesj):

If the present value of $1 for some period and interest rate is .7312, then of course 0.7312 < 1

OpenStudy (jamesj):

In fact, for any rate of interest any time period \( t \geq 1 \), it is the case that \[ 1 > PV(1) = \frac{1}{(1+r)^t} \]

OpenStudy (jamesj):

This is the mathematical expression of a really important fact in finance: a dollar today is worth more than a dollar in the future.